# ThermoHands: A Benchmark for 3D Hand Pose Estimation from Egocentric Thermal Images

### v3

@@ Abstract
In this work, we present ThermoHands, a new benchmark for thermal image-based egocentric 3D hand pose estimation, aimed at overcoming challenges like varying lighting conditions and obstructions (e.g., handwear). The benchmark includes a multi-view and multi-spectral dataset collected from 28 subjects performing hand-object and hand-virtual interactions under diverse scenarios, accurately annotated with 3D hand poses through an automated process. We introduce a new baseline method, TherFormer, utilizing dual transformer modules for effective egocentric 3D hand pose estimation in thermal imagery. Our experimental results highlight TherFormer’s leading performance and affirm thermal imaging’s effectiveness in enabling robust 3D hand pose estimation in adverse conditions.
1 Introduction
@@
Egocentric 3D hand pose estimation is critically important for interpreting hand gestures across various applications, ranging from extended reality (XR) [1, 2, 3, 4], to human-robot interaction [5, 6, 7], and to imitation learning [8, 9, 10]. Its importance has been magnified with the advent of advanced XR headsets such as the Meta Quest series [11] and Apple Vision Pro [12], where it serves as a cornerstone for spatial interaction and immersive digital experiences.
@@
While current research of hand pose estimation primarily focuses on RGB image-based methods [13, 14, 15, 16, 17], these approaches are particularly vulnerable to issues related to lighting variation and occlusions caused by handwear, e.g., gloves or large jewellery [18, 19]. These challenges underscore the imperative for robust egocentric 3D hand pose estimation capable of performing reliably in a variety of common yet complex daily scenarios. The prevailing approach to facilitate robust hand pose estimation in low-light conditions utilizes near infrared (NIR) cameras paired with active NIR emitters. This technology, invisible to the human eye, leverages active NIR emitter-receiver configurations for depth estimation through time-of-flight (ToF) or structured lighting. Nevertheless, active NIR systems are more power-intensive compared to passive sensing technologies [20, 21] and are prone to interference from external NIR sources, such as sunlight [22] and other NIR-equipped devices [23]. Consequently, these vulnerabilities restrict the effectiveness of hand pose estimation under bright daylight conditions and in situations where multiple augmented reality (AR) or virtual reality (VR) systems are used for collaborative works.
@@
In contrast to NIR-based methods, thermal imaging cameras offer a passive sensing solution for hand pose estimation by capturing long-wave infrared (LWIR) radiation emitted from objects, thereby eliminating reliance on the visible light spectrum [24]. This unique attribute of thermal imaging introduces several benefits for hand pose estimation. Primarily, it accentuates the hand’s structure via temperature differentials, negating the effects of lighting variability. Moreover, thermal cameras can detect hands even under handwear such as gloves by identifying heat transmission patterns. This ability ensures a stable and consistent representation of hands, independent of any coverings, thereby broadening the scope and reliability of hand pose estimation across various scenarios.
@@
Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR as two ‘spectra’, despite their usual overlap. spectral imagery? To answer this, we introduce ThermoHands, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging. This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cfFig. 1). Our dataset emulates real-world application contexts by incorporating both hand-object and hand-virtual interaction activities, with participation from 28 subjects to ensure a broad representation of actions. To offer a thorough comparison across spectral types, we gather data under five distinct scenarios, each characterized by varying environments, handwear, and lighting conditions (cfTab. 1). Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline. This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model [25] (cfFig. 3).
@@
Together with the multi-spectral dataset, we introduce a new baseline method named TherFormer, specifically designed for thermal image-based egocentric 3D hand pose estimation (cfFig. 4). This approach is notable for its two consecutive transformer modules, i.e., mask-guided spatial transformer and temporal transformer, which encode spatio-temporal relationship for 3D hand joints without losing the computation efficiency. Our validation process begins with verifying the annotation quality, which averages an accuracy of 1cm (cfTab. 2). We then benchmark TherFormer against leading methods (cfTab. 3) and compare the performance of various spectral images (cfTab. 4 and Fig. 6). The findings underscore thermal imagery’s advantages in difficult lighting conditions and when hands are gloved, showing superior performance and better adaptability to challenging settings than other spectral techniques. Our main contributions are summarized as follows:
@@
2 Related Works
2.1 3D Hand Pose Datasets
Datasets with 3D hand pose annotations are imperative for training and evaluating ad-hoc models. Existing datasets, according to their approaches of annotation acquisition, can be summarized as four types in general, i.e., marker-based [26, 27, 28, 29], synthetic [30, 31, 32, 33, 34], manual [30, 35] or hybrid [36, 37, 38, 39, 40], and automatic [41, 42, 43, 44] annotated datasets. Marker-based approaches, using magnetic sensor [26, 27] or Mocap markers [28, 29], can alter and induce bias to the hand appearance. Synthetic data [30, 31, 33, 34, 32] suffers from the sim2real gap in terms of hand motion and texture features. Introducing human annotators to fully [30, 35] or partly [36, 37, 38, 39, 40] annotate 2D/3D keypoints circumvents the issues above, but it either limits the scale of datasets or manifests costly and laborious in practice. Most similar to ours, some datasets adopt fully automatic pipelines to obtain 3D hand pose annotations [41, 42, 43], which leverage pre-trained models (e.g. OpenPose [45]) to infer the prior hand information and rely on optimization to fit the MANO hand model [25].
Despite the existing progress, previous datasets only provide depth [26, 32], RGB images [33, 38, 37, 40] or both of them [27, 28, 31, 34, 30, 35, 36, 39, 41, 42, 43] as the input spectra, unable to support the study of NIR or thermal image-based 3D hand pose estimation. ThermoHands fills the gap by providing a moderate amount of multi-spectral image data, from infrared to visual light, paired with depth images. Moreover, we capture bimanual actions from both egocentric and exocentric viewpoints and design hand-object as well as hand-virtual interaction actions to facilitate a wide range of applications.
Figure 1: Data capture setup with the HMSP and exocentric platform recording multi-view multi-spectral images of two-hand actions performed by participants.
2.2 Image-based 3D Hand Pose Estimation
As a key computer vision task, 3D hand pose estimation from images is highly demanded by applications like XR [1, 2, 3, 4], human-robot interaction [5, 6, 7] and imitation learning [8, 9, 10]. Therefore, this field has been extensively explored in previous arts that uses single RGB [46, 47, 48, 49, 50, 13, 38, 14, 51, 52, 53, 31, 54, 55, 56, 57] or depth [58, 59, 60, 61, 62] image as input. These methods can be roughly categorized into two fashions, i.e., model-based and model-free methods. Model-based methods [14, 51, 52, 53, 31, 54, 55, 56, 57] utilize the prior knowledge of the MANO hand model [25] by estimating its shape and pose parameters, while model-free methods [46, 47, 48, 49, 50, 13, 38, 58, 59, 60, 61, 62] learn the direct regression of 3D hand joints or vertices coordinates. Recently there has been growing interest in leveraging the temporal supervision [63, 54, 64, 65, 66] or leveraging sequential images as input [15, 16, 17, 67, 68, 69] for 3D hand pose estimation. In this study, we evaluate existing methods and our baseline method in both single image-based and video-based problem settings, respectively. Apart from the previous approaches, we investigate the potential of thermal imagery for tackling various challenges in 3D hand pose estimation.
2.3 Thermal Computer Vision
Thermal cameras achieve imaging by capturing the radiation emitted in the LWIR spectrum and deducing the temperature distribution on the surfaces [24]. Leveraging its robustness to variable illumination and unique temperature information, numerous efforts have been made to address various computer vision tasks, including super-resolution [70, 71, 72], human detection [73, 74, 75], action recognition [76, 77] and pose estimation [78, 79, 80], semantic segmentation [81, 82, 83, 84], depth estimation [85, 86, 87, 88], visual(-inertial) odometry/SLAM [89, 90, 91, 92], 3D reconstruction [93, 94, 95], etc. In this work, we focus on 3D hand pose estimation, which is an under-exploited task based on thermal images.
3 The ThermoHands Benchmark
3.1 Multi-Spectral Hand Pose Dataset
Overview. At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cfTab. 1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^2] The study has received the ethical approval from School of Informatics, University of Edinburgh, and participant consent forms were signed before the collection.. As shown in Fig. 1, we develop a customized head-mounted sensor platform (HMSP) and an exocentric platform to record multi-view data. During the capture, our participants are asked to perform pre-defined hand-object and hand-virtual interaction actions within the playground above the table. The main part is captured in the normal office scenario. To facilitate the evaluation under different settings, four auxiliary parts are recorded i) under the darkness, ii) under the sun glare, iii) with gloves on hand, and iv) in the kitchen environment with different actions, respectively.
@@ 3.1 Multi-Spectral Hand Pose Dataset
Table 1: Benchmark Dataset Statistics. The overall duration of our dataset is over 3 hours with $\sim$96K synchronized frame of all types of images collected.
@@
Figure 2: Thermal calibration chessboard containing a black base board and multiple removable white cubes (a). By cooling down the base board, it shows similar patterns and allows automatic corner detection in all (b) RGB, (c) NIR and (d) thermal images.
@@
Sensor Platforms. The HMSP is mounted with an Intel RealSense L515 LiDAR depth camera [20] streaming egocentric RGB, depth and NIR images, and a FLIR Boson 640 thermal camera [96] receiving the LWIR to obtain the thermal images. To better simulate XR devices, the HMSP is modularly designed as two components: a base and a sensor board with a fixed 30-degree downward tilt. An extra exocentric platform equipped with an Intel RealSense D455 [21] is leveraged to support multi-view annotation (cf. Sec. 3.2) as well as provide the RGB-D image data from the third-person viewpoint. As exhibited in Fig. 1, we place the two depth sensors outside each other’s field of view (FoV) to minimize interference caused by their NIR emitters [20, 21]. In the office environment, random heat sources, e.g., servers and chargers, are strategically placed in the background to increase realism and introduce challenging factors into thermal images.
@@
Calibration. For accuracy, we use factory-calibrated intrinsic parameters from the D455 and L515 cameras [20, 21]. To better calibrate the parameters of the thermal camera, we self-design a modular calibration chessboard as shown in Fig. 2. Before calibration, we cool down the black base board while keeping the white cubes at room temperature to create a visible chessboard pattern in all three spectra, which can be seen in Fig. 2. In this way, we can simultaneously calibrate the intrinsic parameter of the L515 thermal camera and its extrinsic parameter w.r.t. the D455 RGB-D camera. To enable the calibration between the two viewpoints, we place 11 IR reflective spheres at random locations and heights above the table. At the first frame of each sequence, we manually annotate their 2D locations from two viewpoints, retrieve their depth values and compute the transformation between two viewpoints by solving the PnP problem [97]. For the subsequent frames, as the exocentric platform keeps stationary, we only track the the motion of the ego-head using KISS-ICP [98] odometry method, which takes the point clouds converted from the egocentric depth image as input.
@@
Dataset Statistics. As seen in Tab. 1, our dataset consists of approximately 96K synchronized multi-spectral multi-view frames (cfFig. 1) and 352 independent sequences in total. The main part is collected under the normal office scenario, where each participant [^3] Due to their limited time, 7 participants only perform the hand-object actions. performs 7 hand-object interaction actions: cut paper, fold paper, pour water, read book, staple paper, write with pen, write with pencil, and 5 hand-virtual interaction actions: pinch and drag, pinch and hold, swipe, tap, touch, with two hands. This main part is divided into the training, validation and testing splits by subjects with a ratio of 4:1:2. We also collect four auxiliary testing sets by asking one subject to perform the aforementioned 12 actions in the darkness, sun glare and gloves settings individually, and two subjects to perform 7 scenarios-specific interaction actions: cut, spray, stir, wash hands, wash mug, wash plate, wipe in the kitchen. Please refer to the supplementary for more dataset details.
3.2 Hand Pose Annotation
To avoid employing tedious human efforts for annotation, we implement a fully automatic annotation pipeline, similar to the approaches in [41, 42, 43], to obtain the 3D hand pose ground truth for our dataset. In particular, we use the MANO statistical hand model [25] to represent 3D hand pose. The MANO model parameterizes the hand mesh vertices into two low-dimensional embeddings, i.e., the shape parameters $\beta\in\mathbb{R}^{10}$ and the pose parameters $\theta\in\mathbb{R}^{51}$. The MANO fitting is performed by minimizing the following optimization objective per frame for each hand:
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\lambda_{j2d}\mathcal{L}_{j2d}^{{\color[rgb]{1,.5,0}\bullet}}+\lambda_{mask}\mathcal{L}_{mask}^{{\color[rgb]{1,0,1}\bullet}}+\lambda_{j3d}\mathcal{L}_{j3d}^{{\color[rgb]{0,1,0}\bullet}}+\lambda_{mesh}\mathcal{L}_{mesh}^{{\color[rgb]{0,1,1}\bullet}}+\lambda_{reg}\mathcal{L}_{reg}^{{\color[rgb]{1,0,0}\bullet}}$ (1)
where $\lambda_{j2d},\lambda_{mask},\lambda_{j3d},\lambda_{mesh},\lambda_{reg}$ are used to balance the weight of different errors. The diagram illustrating our annotation process is shown in Fig. 3. For more explanation of error terms and implementation details of our dataset annotation process, please refer to our supplementary document.

### v4

@@ Abstract.
Designing egocentric 3D hand pose estimation systems that can perform reliably in complex, real-world scenarios is crucial for downstream applications. Previous approaches using RGB or NIR imagery struggle in challenging conditions: RGB methods are susceptible to lighting variations and obstructions like handwear, while NIR techniques can be disrupted by sunlight or interference from other NIR-equipped devices. To address these limitations, we present ThermoHands, the first benchmark focused on thermal image-based egocentric 3D hand pose estimation, demonstrating the potential of thermal imaging to achieve robust performance under these conditions. The benchmark includes a multi-view and multi-spectral dataset collected from 28 subjects performing hand-object and hand-virtual interactions under diverse scenarios, accurately annotated with 3D hand poses through an automated process. We introduce a new baseline method, TherFormer, utilizing dual transformer modules for effective egocentric 3D hand pose estimation in thermal imagery. Our experimental results highlight TherFormer’s leading performance and affirm thermal imaging’s effectiveness in enabling robust 3D hand pose estimation in adverse conditions.
* Equal contribution
† Corresponding author: xiaoxuan.lu@ucl.ac.uk
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: none
Figure 1. Data capture setup with the customized head-mounted sensor platform (HMSP) and exocentric platform recording multi-view multi-spectral images of two-hand actions performed by participants.
1. Introduction
@@
Egocentric 3D hand pose estimation is critically important for interpreting hand gestures across various applications, ranging from extended reality (XR) (wu2020hand, ; sagayam2017hand, ; marchand2015pose, ; liang2015ar, ), to human-robot interaction (sampieri2022pose, ; gao2021dynamic, ; gao2019dual, ), and to imitation learning (qin2022dexmv, ; chen2023bi, ; antotsiou2018task, ). Its importance has been magnified with the advent of advanced XR headsets such as the Meta Quest series (meta_quest, ) and Apple Vision Pro (apple_vision_pro_2024, ), where it serves as a cornerstone for spatial interaction and immersive digital experiences.
@@
While current research of hand pose estimation primarily focuses on RGB image-based methods (lin2021two, ; li2022interacting, ; wen2023hierarchical, ; Fu_2023_ICCV, ; cho2023transformer, ), these approaches are particularly vulnerable to issues related to lighting variation and occlusions caused by handwear, e.g., gloves or large jewellery (youtube_video_2024, ; apple_support_gloves_2024, ). These challenges underscore the imperative for robust egocentric 3D hand pose estimation capable of performing reliably in a variety of common yet complex daily scenarios. The prevailing approach to facilitate robust hand pose estimation in low-light conditions utilizes near infrared (NIR) cameras paired with active NIR emitters. This technology, invisible to the human eye, leverages active NIR emitter-receiver configurations for depth estimation through time-of-flight (ToF) or structured lighting. Nevertheless, active NIR systems are more power-intensive compared to passive sensing technologies (IntelRSL515, ; IntelRealSense2023D455, ) and are prone to interference from external NIR sources, such as sunlight (suarez2012using, ) and other NIR-equipped devices (seewald2019toward, ). Consequently, these vulnerabilities restrict the effectiveness of hand pose estimation under bright daylight conditions and in situations where multiple augmented reality (AR) or virtual reality (VR) systems are used for collaborative works.
@@
In contrast to NIR-based methods, thermal imaging cameras offer a passive sensing solution for hand pose estimation by capturing long-wave infrared (LWIR) radiation emitted from objects, thereby eliminating reliance on the visible light spectrum (lloyd2013thermal, ). This unique attribute of thermal imaging introduces several benefits for hand pose estimation. Primarily, it accentuates the hand’s structure via temperature differentials, negating the effects of lighting variability. Moreover, thermal cameras can detect hands even under handwear such as gloves by identifying heat transmission patterns. This ability ensures a stable and consistent representation of hands, independent of any coverings, thereby broadening the scope and reliability of hand pose estimation across various scenarios.
@@
Building on the above insights, this study probes the following research question: Can egocentric thermal imagery be effectively used for 3D hand pose estimation under various conditions (such as different lighting and handwear), and how does it compare to techniques using RGB, NIR, and depth [^1] For readability, we treat depth and NIR images as two ‘spectra’, despite their usual overlap. spectral imagery? To answer this, we introduce ThermoHands  [^2] Anonymous project page: https://thermohands.github.io/, the first benchmark specifically tailored for egocentric 3D hand pose estimation utilizing thermal imaging. This benchmark is supported by a novel multi-spectral and multi-view dataset designed for egocentric 3D hand pose estimation and is unique in comprising thermal, NIR, depth, and RGB images (cfFig. 1). Our dataset emulates real-world application contexts by incorporating both hand-object and hand-virtual interaction activities, with participation from 28 subjects to ensure a broad representation of actions. To offer a thorough comparison across spectral types, we gather data under five distinct scenarios, each characterized by varying environments, handwear, and lighting conditions (cfTab. 1). Considering the challenges associated with manually annotating large-scale 3D hand poses, we developed an automated annotation pipeline. This pipeline leverages multi-view RGB and depth imagery to accurately and efficiently generate 3D hand pose ground truths through optimization based on the MANO model (MANO, ) (cfFig. 4).
@@
Together with the multi-spectral dataset, we introduce a new baseline method named TherFormer, specifically designed for thermal image-based egocentric 3D hand pose estimation (cfFig. 5). This approach is notable for its two consecutive transformer modules, i.e., mask-guided spatial transformer and temporal transformer, which encode spatio-temporal relationship for 3D hand joints without losing the computation efficiency.
Our validation process begins with verifying the annotation quality, which averages an accuracy of 1cm (cfTab. 2). We then benchmark TherFormer against leading methods (cfTab. 3) and compare the performance of various spectral images (cfTab. 5, Tab. 4 and Fig. 8). The findings underscore thermal imagery’s advantages in difficult lighting conditions and when hands are gloved, showing superior performance and better adaptability to challenging settings than other spectral techniques. Our main contributions are summarized as follows:
@@
2. Related Works
@@
Table 1. Benchmark Dataset Statistics. The overall duration of our dataset is over 3 hours with $\sim$96K synchronized frame of all types of images collected.
2.1. 3D Hand Pose Datasets
Datasets with 3D hand pose annotations are imperative for training and evaluating ad-hoc models. Existing datasets, according to their approaches of annotation acquisition, can be summarized as four types in general, i.e., marker-based (yuan2017bighand2, ; garcia2018first, ; taheri2020grab, ; fan2023arctic, ), synthetic (mueller2017real, ; hasson2019learning, ; mueller2019real, ; mueller2018ganerated, ; zimmermann2017learning, ), manual (mueller2017real, ; sridhar2016real, ) or hybrid (chao2021dexycb, ; zimmermann2019freihand, ; moon2020interhand2, ; liu2022hoi4d, ; ohkawa2023assemblyhands, ), and automatic (hampali2020honnotate, ; brahmbhatt2020contactpose, ; kwon2021h2o, ; sener2022assembly101, ) annotated datasets. Marker-based approaches, using magnetic sensor (yuan2017bighand2, ; garcia2018first, ) or Mocap markers (taheri2020grab, ; fan2023arctic, ), can alter and induce bias to the hand appearance. Synthetic data (mueller2017real, ; hasson2019learning, ; mueller2018ganerated, ; zimmermann2017learning, ; mueller2019real, ) suffers from the sim2real gap in terms of hand motion and texture features. Introducing human annotators to fully (mueller2017real, ; sridhar2016real, ) or partly (chao2021dexycb, ; zimmermann2019freihand, ; moon2020interhand2, ; liu2022hoi4d, ; ohkawa2023assemblyhands, ) annotate 2D/3D keypoints circumvents the issues above, but it either limits the scale of datasets or manifests costly and laborious in practice. Most similar to ours, some datasets adopt fully automatic pipelines to obtain 3D hand pose annotations (hampali2020honnotate, ; brahmbhatt2020contactpose, ; kwon2021h2o, ), which leverage pre-trained models (e.g. OpenPose (8765346, )) to infer the prior hand information and rely on optimization to fit the MANO hand model (MANO, ).
Despite the existing progress, previous datasets only provide depth (yuan2017bighand2, ; mueller2019real, ), RGB images (mueller2018ganerated, ; moon2020interhand2, ; zimmermann2019freihand, ; ohkawa2023assemblyhands, ) or both of them (garcia2018first, ; taheri2020grab, ; hasson2019learning, ; zimmermann2017learning, ; mueller2017real, ; sridhar2016real, ; chao2021dexycb, ; liu2022hoi4d, ; hampali2020honnotate, ; brahmbhatt2020contactpose, ; kwon2021h2o, ) as the input spectra, unable to support the study of NIR or thermal image-based 3D hand pose estimation. ThermoHands fills the gap by providing a moderate amount of multi-spectral image data, from infrared to visual light, paired with depth images. Moreover, we capture bimanual actions from both egocentric and exocentric viewpoints and design hand-object as well as hand-virtual interaction actions to facilitate a wide range of applications.
2.2. Image-based 3D Hand Pose Estimation
As a key computer vision task, 3D hand pose estimation from images is highly demanded by applications like XR (wu2020hand, ; sagayam2017hand, ; marchand2015pose, ; liang2015ar, ), human-robot interaction (sampieri2022pose, ; gao2021dynamic, ; gao2019dual, ) and imitation learning (qin2022dexmv, ; chen2023bi, ; antotsiou2018task, ). Therefore, this field has been extensively explored in previous arts that uses single RGB (jiang2023a2j, ; spurr2020weakly, ; yang2019disentangling, ; cheng2021handfoldingnet, ; kim2021end, ; lin2021two, ; moon2020interhand2, ; li2022interacting, ; smith2020constraining, ; wang2020rgb2hands, ; zhang2021interacting, ; hasson2019learning, ; liu2021semi, ; boukhayma20193d, ; baek2019pushing, ; zhang2019end, ) or depth (ge2018robust, ; moon2018v2v, ; xiong2019a2j, ; zhang2020handaugment, ; zhu2019hand, ) image as input. These methods can be roughly categorized into two fashions, i.e., model-based and model-free methods. Model-based methods (li2022interacting, ; smith2020constraining, ; wang2020rgb2hands, ; zhang2021interacting, ; hasson2019learning, ; liu2021semi, ; boukhayma20193d, ; baek2019pushing, ; zhang2019end, ) utilize the prior knowledge of the MANO hand model (MANO, ) by estimating its shape and pose parameters, while model-free methods (jiang2023a2j, ; spurr2020weakly, ; yang2019disentangling, ; cheng2021handfoldingnet, ; kim2021end, ; lin2021two, ; moon2020interhand2, ; ge2018robust, ; moon2018v2v, ; xiong2019a2j, ; zhang2020handaugment, ; zhu2019hand, ) learn the direct regression of 3D hand joints or vertices coordinates. Recently there has been growing interest in leveraging the temporal supervision (hasson2020leveraging, ; liu2021semi, ; yang2020seqhand, ; Kocabas_2020_CVPR, ; park2022handoccnet, ) or leveraging sequential images as input (wen2023hierarchical, ; Fu_2023_ICCV, ; cho2023transformer, ; Cai_2019_ICCV, ; khaleghi2022multi, ; choi2021beyond, ) for 3D hand pose estimation. In this study, we evaluate existing methods and our baseline method in both single image-based and video-based problem settings, respectively. Apart from the previous approaches, we investigate the potential of thermal imagery for tackling various challenges in 3D hand pose estimation.
Figure 2. Design of the head-mounted sensor platform and sensor alignment.
2.3. Thermal Computer Vision
Thermal cameras achieve imaging by capturing the radiation emitted in the LWIR spectrum and deducing the temperature distribution on the surfaces (lloyd2013thermal, ). Leveraging its robustness to variable illumination and unique temperature information, numerous efforts have been made to address various computer vision tasks, including super-resolution (rivadeneira2023thermal, ; gupta2021toward, ; kansal2020multi, ), human detection (10.5555/3451271.3451272, ; ivavsic2019human, ; haider2021human, ), action recognition (batchuluun2019action, ; ding2022individual, ) and pose estimation (lupion20243d, ; chen2020multi, ; smith2023human, ), semantic segmentation (kutuk2022semantic, ; kim2021ms, ; vertens2020heatnet, ; li2020segmenting, ), depth estimation (Shin_2023_CVPR, ; lu2021alternative, ; shin2021self, ; kim2018multispectral, ), visual(-inertial) odometry/SLAM (rob.21932, ; 9623261, ; shin2019sparse, ; saputra2020deeptio, ), 3D reconstruction (liu2023humans, ; sage20213d, ; schramm2022combining, ), etc. In this work, we focus on 3D hand pose estimation, which is an under-exploited task based on thermal images.
@@ 2.3. Thermal Computer Vision
Figure 3. Thermal calibration chessboard containing a black base board and multiple removable white cubes (a). By cooling down the base board, it shows similar patterns and allows automatic corner detection in all (b) RGB, (c) NIR and (d) thermal images.
3. The ThermoHands Benchmark
3.1. Multi-Spectral Hand Pose Dataset
Overview. At the core of our benchmark lies a multi-spectral dataset for 3D hand pose estimation (cfTab. 1), capturing hand actions performed by 28 subjects of various ethnicities and genders [^3] The study has received the ethical approval from XXX (no name for blind review), and participant consent forms were signed before the collection.. As shown in Fig. 1, we develop a customized head-mounted sensor platform (HMSP) and an exocentric platform to record multi-view data. During the capture, our participants are asked to perform pre-defined hand-object and hand-virtual interaction actions within the playground above the table. The main part is captured in the normal office scenario. To facilitate the evaluation under different settings, four auxiliary parts are recorded i) under the darkness, ii) under the sun glare, iii) with gloves on hand, and iv) in the kitchen environment with different actions, respectively.
@@ 3.1. Multi-Spectral Hand Pose Dataset
Sensor Platforms. Apart from the traditional approach of mounting cameras on helmets (kwon2021h2o, ), the design of the HMSP (cfFig. 2) focuses on simulating an actual Mixed Reality (MR) device and reducing extra weight to allow participants to perform freely. The HMSP consists of three major components: a cushion for comfort, a base component that provides a 30-degree downward tilt, and a sensor board that carries two cameras - an Intel RealSense L515 LiDAR camera (IntelRSL515, ) that streams egocentric RGB, depth, and NIR images, and a Teledyne FLIR Boson 640 long-wave infrared (LWIR) camera (FLIRBoson2024, ) that receives the LWIR to obtain the thermal images. An extra exocentric platform equipped with an Intel RealSense D455 (IntelRealSense2023D455, ) is leveraged to support multi-view annotation (cf. Sec. 4) as well as provide the RGB-D image data from the third-person viewpoint. As exhibited in Fig. 1, we place the two depth sensors outside each other’s field of view (FoV) to minimize interference caused by their NIR emitters (IntelRSL515, ; IntelRealSense2023D455, ).
@@
Egocentric Calibration. For accuracy, factory-calibrated intrinsic parameters are used for the Intel RealSense D455 and L515 cameras (IntelRSL515, ; IntelRealSense2023D455, ). As seen in Fig. 2, although the thermal camera’s center is aligned with the RGB camera’s center at the same height on the 3D-printed sensor board, we still calibrate the extrinsic parameters between these two egocentric cameras to ensure enhanced alignment accuracy. To accommodate the unique imaging mechanism of thermal camera, we self-design a modular calibration chessboard as shown in Fig. 3. Before calibration, we cool down the black base board while keeping the white cubes at room temperature to create a visible chessboard pattern in all three spectra, which can be seen in Fig. 3. In this way, we can simultaneously calibrate the intrinsic parameter of the thermal camera and its extrinsic parameter w.r.t. the D455 RGB-D camera.
Cross-View Calibration. To calibrate the cameras between the egocentric and exocentric viewpoints, we place 11 IR reflective spheres at random heights within the playground, ensuring they are visible from both viewpoints. Initially, we plan to detect these spheres automatically from the NIR images and track them with the Kalman Filter (wan2000unscented, ). However, we find this approach leads to many false positive detections, which severely affect the tracking accuracy. To ensure the calibration accuracy, we only annotate the sphere markers manually from two viewpoints in the first frame. The 3D positions of markers can be computed using the egocentric depth image and the cross-view transformation can be computed by solving the PnP (FischlerBolles1981, ). For the subsequent frames, we can calculate the transformation by only tracking the pose of the egocentric camera as the exocentric platform keeps stationary during collection. We leverage the state-of-the-art odometry method KISS-ICP (vizzo2023kiss, ) and track the motion of the egocentric camera with the point clouds converted from depth images. Specifically, we set the initial_threshold and min_motion_th parameters of KISS-ICP as 0.0001 to make it capable of catching subtle motion. The input point cloud range is set as [0.2, 2.0] meters while the max_points_per_voxel is 30.
@@
Dataset Statistics. As seen in Tab. 1, our dataset consists of approximately 96K synchronized multi-spectral multi-view frames (cfFig. 1) and 352 independent sequences in total. The main part is collected under the normal office scenario, where each participant [^4] Due to their limited time, 7 participants only perform the hand-object actions. performs 7 hand-object interaction actions: cut paper, fold paper, pour water, read book, staple paper, write with pen, write with pencil, and 5 hand-virtual interaction actions: pinch and drag, pinch and hold, swipe, tap, touch, with two hands. This main part is divided into the training, validation and testing splits by subjects with a ratio of 4:1:2. We also collect four auxiliary testing sets by asking one subject to perform the aforementioned 12 actions in the darkness, sun glare and gloves settings individually, and two subjects to perform 7 scenarios-specific interaction actions: cut, spray, stir, wash hands, wash mug, wash plate, wipe in the kitchen.